MN blood type in humans is determined by two alleles at a single co-dominant locus; these alleles are LM and LN. In a sample of 1190 individuals , it was found that:

75 had type M blood (LMLM)

800 had type MN blood (LMLN)

315 had type N blood (LNLN)

Given this population data, calculate the observed allele frequencies and the expected number of individuals of each genotype. For allele frequency reposes, please round to 3 decimal places (ex. 0.000). For individuals, please round to the nearest integer (ex. 123).What is the observed frequency of the LM allele? What is the observed frequency of the LN allele? Given the assumptions of th
Answer #1

1

LMLM =75

LMLN=800

LNLN = 315

total = 1190

frequency of Lm = 2* LMLM + LMLN/ 2* total alleles

= 2*75+800/2*1190

= 150.33 of individuals expected to have blood group M

frequency of Ln = 2*LNLN+ LMLN/2*1190

= 2*315 + 800/ 2*1190

= 630.33 of individuals expected to have blood group N

frequncy of Lmn = 1190 - (150.33+630.33)

= 409.34 of individuals expected to have blood group MN